# On the relationship between Koopman operator approximations and neural ordinary differential equations for data-driven time-evolution predictions

### v1

@@ III.5 Continuous-time models from EDMD
As noted in Sec. II, the Koopman formalism also gives rise to a continuous-time representation of the dynamics on the observable space, described by the infinitesimal generator, $\mathcal{L}$. Using this formulation, an analagous projection step can be performed, which directly yields a neural ODE representation of the state space dynamics, as illustrated in Fig. 2. Based on this, several previous works have shown that continuous-time representations can also be derived from EDMD or similar regression methods [62, 63, 64, 65, 66]. Similarly to standard EDMD, these methods produce a matrix representation $\mathbf{L}$ of the Koopman generator, $\mathcal{L}$ , which acts on observables $g_{\mathbf{c}}=\mathbf{c}^{\mathrm{T}}\Psi\in\mathcal{O}_{\mathbb{D}}$ by updating the projection coefficients:
$\mathcal{L}g_{\mathbf{c}}=(\mathbf{L}\mathbf{c})^{\mathrm{T}}\Psi=\mathbf{c}^{\mathrm{T}}\mathbf{L}^{\mathrm{T}}\Psi+r_{\mathbf{c}}(\mathbf{x})$ (30)
@@
where, as before, the residual $r_{\mathbf{c}}(\mathbf{x})$ arises due to the space $\mathcal{O}_{\mathbb{D}}$ not being Koopman invariant. In many implementations, it is assumed that time derivative data of the states $\dot{\mathbf{x}}$ is available (or can be obtained by finite-differencing). If this velocity data is available, then using the generator definition in Eq. 6 we have
$\mathcal{L}g_{\mathbf{c}}=\frac{\partial g_{\mathbf{c}}}{\partial\mathbf{x}}\cdot\dot{\mathbf{x}}=\mathbf{c}\cdot\frac{\partial\Psi(\mathbf{x})}{\partial\mathbf{x}}\cdot\dot{\mathbf{x}}\,.$ (31)
Then equating Eqs. 30 and 31, we see that the residual can be written as
$r_{\mathbf{c}}(\mathbf{x})=\mathbf{c}^{\mathrm{T}}\left(\frac{\partial\Psi(\mathbf{x})}{\partial\mathbf{x}}\cdot\dot{\mathbf{x}}-\mathbf{L}^{\mathrm{T}}\Psi\right)$ (32)
and so we can frame the EDMD regression for $\mathbf{L}$ by minimizing the objective
@@
where $\dot{\Psi}(\mathbf{x})=\frac{\partial\Psi}{\partial\mathbf{x}}\cdot\dot{\mathbf{x}}$. This is the approach taken in [64, 65, 66]. However, this is disadvantageous, as time derivative data can be difficult to acquire from experiments, and obtaining accurate finite-differencing approximations relies on measurements being closely spaced in time. Alternatively, one can obtain a matrix approximation of the generator from the EDMD approximation of the Koopman operator by applying the limit definition in Eq. 5 as
@@
though this approach will only be accurate for sufficiently small time interval, $\tau$. Similarly, we can make use of the exponential relationship between the Koopman operator and its generator in Eq. 7 and obtain the matrix approximation of $\mathcal{L}$ using a matrix logarithm as
@@
Once the matrix representation of the generator is obtained, it can be used to predict the dynamics of the state, $\mathbf{x}$, as this matrix acts on the dictionary elements to return their time derivative. Specifically, given an initial state $\mathbf{x}_{0}$, we can lift it to obtain an initial condition for the lifted state, $\Psi_{0}=\Psi(\mathbf{x}_{0})$, and then obtain a trajectory in this lifted state by integrating the equation
@@
Integrating this yields a trajectory in the lifted state, $\Psi$, from which we can recover the state predictions by mapping back to the state space, $\mathbf{x}=\mathcal{P}\Psi$, as in Sec. III.4. Note that Eq. 36 is a linear ODE in $\Psi$ and is essentially a continuous time version of the linear time evolution using the Koopman operator (without projection) in Eqs. 22, 25.
Analogously the EDMD formulation, where we defined a nonlinear discrete-time evolution equation on the state space using the EDMD operator, $\mathbf{K}_{\tau}$, by projecting back to the state space on each timestep (see Eq. 26), we can also define a nonlinear continuous time evolution equation on the state space using the generator obtained from EDMD by lifting the state $\Psi(\mathbf{x})$, computing the time derivative of the feature vector, and then projecting the time derivative back to the state space. This can be seen by simply taking the time derivative of the projection back to the state space $\mathbf{x}=\mathcal{P}\Psi(\mathbf{x})$ as
@@
This form is only necessary if the mapping $\mathcal{P}$ back to the state space is nonlinear (such as in Koopman autoencoder formulations). If $\mathcal{P}$ is taken to be a linear mapping, $\mathcal{P}\Psi=\mathbf{P}\Psi$, such as when the state is explicitly included in the dictionary (see Sec. III.4, Eq. 23), then $\frac{\partial\mathcal{P}}{\partial\Psi}=\mathbf{P}$ and Eq. 37 can be simplified to
@@
In the case of EDMD-DL, where the feature vector $\Psi(\mathbf{x})$ is parametrized by a neural network, $\Psi(\mathbf{x};\theta_{\Psi})$ and optimized jointly with $\mathbf{K}_{\tau}$, this implies that the EDMD-DL training procedure directly defines a neural ODE of a certain structure
@@
In summary, in both discrete time and continuous time, performing a projection back to the state space in the EDMD time evolution leads to a nonlinear dynamical system in the state space with a natural structure for neural networks. Specifically, the state is first expanded nonlinearly into a high dimensional feature space, followed by a linear mapping which represents the dynamics as a linear combination of these features. In the following section, we see how this observation motivates the development of several methods that combine different aspects of EDMD and neural ODE based models. We also see that the projection within EDMD-DL time evolution leads to greatly improved performance as compared to the linear time evolution strategy used in standard EDMD-DL. Additionally, the models using this projection achieve a level of accuracy in predicting extreme events comparable to the non-Markovian approach of Racca and Magri [67].
IV Performance comparison
IV.1 Model structures
The discussion in Sec. III of the parallels between data-driven state-space modeling techniques and the projected Koopman approach to predicting the state evolution motivates a comparison between the numerical methods arising from each. In particular, we will compare two neural ODE models arising directly from the state space view and four neural network models arising from EDMD approximations of the Koopman operator (with learned dictionaries). These model variations are described in the text below and summarized in Table 1.
Table 1: Summary of model structures based on neural ODEs and EDMD-DL
@@ IV.1 Model structures
EDMD structured neural ODE $\dot{\mathbf{x}}=\mathbf{P}\mathbf{L}^{\mathrm{T}}\Psi(\mathbf{x};\theta_{\Psi})$ $\theta_{\Psi}$ and matrix $\mathbf{P}\mathbf{L}^{\mathrm{T}}$ optimized
@@
EDMD basis neural ODE $\dot{\mathbf{x}}=\mathbf{P}\mathbf{L}^{\mathrm{T}}\Psi(\mathbf{x};\theta_{\Psi})$ $\theta_{\Psi}$ optimized by EDMD-DL (Eq. 24)
$\mathbf{x}(0)=\mathbf{x}_{0}$ $\mathbf{P}\mathbf{L}^{\mathrm{T}}$ optimized as neural ODE (Eq. 11)
@@
The first EDMD approach that we consider is the standard EDMD-DL method, where a dictionary paremetrized by a neural network and the corresponding Koopman operator matrix representation are jointly learned by optimizing the objective in Eq. 24. Then to model the time-evolution, the initial state is lifted and propagated linearly in the lifted space using the Koopman matrix (see Eq. 25). We also consider the EDMD variation with reprojection, where the same dictionary and Koopman matrix are used, but where on each timestep, the state is lifted, evolved forward in time, and then reprojected (see. Eq. 26). In each EDMD-DL implementation, we constrain the dictionary to be of the form in Eq. 23, where the state is explicitly included as a nontrainable set of dictionary elements.
@@
We consider two variations of continuous-time models derived from EDMD. In the first, we use the dictionary obtained from EDMD-DL as the neural network in a neural ODE. That is, the dictionary is optimized, as in EDMD-DL, to minimize Eq. 24, and then the linear layer that maps to the output is optimized through the neural ODE training procedure. So the model has the structure of Eq. 39, where the dictionary $\Psi(\mathbf{x};\theta_{\Psi})$ is determined by EDMD-DL and the linear mapping $\mathbf{P}\mathbf{L}^{\mathrm{T}}$ is trained as with the neural ODE loss in Eq. 11. We refer to this variation as an EDMD basis neural ODE. Lastly, we consider a neural ODE model that is obtained directly from EDMD-DL, without a separate training procedure. That is, the matrix representation of the generator $\mathbf{L}$ is obtained from the EDMD-DL Koopman matrix $\mathbf{K}_{\tau}$ by a matrix logarithm, as in Eq. 35. We refer to this variation as an EDMD direct neural ODE.
@@ IV.2 Lorenz system
Finally, while Figs. 3 and 4 analyzed the short-time performance of these models, we demonstrate the long-time performance of the models by simulating each model for $100$ time units from a common initial condition. The resulting trajectory from each model is shown in Fig. 5. We see that each of the models considered (excluding the standard EDMD-DL approach) are able to capture the dynamics and reproduce the butterfly shape of the Lorenz attractor.
@@ IV.3 Low dimensional model of turbulent shear flow
The domain is wall-bounded at $y=\pm 1$ and periodic in the streamwise and spanwise directions ($x$ and $z$), with lengths $L_{x}=4\pi$ and $L_{z}=2\pi$, respectively; and we consider a channel Reynolds number of $\mathrm{Re}=400$, following Refs. [71, 67, 28, 26]. The Lyapunov timescale associated with these parameter values is $\tau_{L}\approx 61$ [67]. The nine mode model is constructed as a Galerkin projection of the Navier-Stokes equations onto nine Fourier modes, as described in [68]. That is, the fluid velocity field $\mathbf{u}(\mathbf{x},t)$ is approximated as a superposition of the modes, $\mathbf{u}_{i}(\mathbf{x})$, with time-varying amplitudes $a_{i}(t)$
@@
Performing the Galerkin projection yields a set of nine ODEs for the amplitudes. These ODEs, as well as the Fourier modes for the model, are given explicitly in Ref. [68]. The first mode of the model, $\mathbf{u}_{1}(\mathbf{x})$, represents the base flow profile, so the laminar flow state is associated with the amplitudes $\mathbf{a}=[1,0,0,0,0,0,0,0,0]^{\mathrm{T}}$. The flow can be characterized using the energy $E(t)=\sum_{i=1}^{9}a_{i}^{2}(t)$. In the turbulent region of the state space, the system shows large bursts of energy, corresponding to quasi-laminarization events, while the laminar state corresponds to the steady value of $E=1$. A typical time-series of the energy for this system is shown in Fig. 6
Refer to captionExtreme eventsLaminarization
@@
A training dataset is constructed by numerically integrating the system from 100 random initial conditions, $\mathbf{a}(0)\sim\mathcal{U}(-0.1,0.1)$ for 30,000 time units or until laminarization occurs with data stored at time intervals of $\tau=0.5$. The first $500$ time units of each trajectory are discarded as transience. These trajectories have a mean lifetime of approximately $1.1\times 10^{4}$ (or $180\,\tau_{L}$) and the resulting dataset contains approximately $2.1\times 10^{6}$ snapshots. A test dataset is generated using an identical procedure, using another set of randomly selected initial conditions. All models are trained for 10 epochs with batch size of 40 ($5.25\times 10^{5}$ iterations).
@@
Refer to caption(a)(b)0.67(c)1.14(d)0.57(e)3.83(f)1.65
Figure 8: Joint probability densities of amplitudes $a_{1}$ and $a_{3}$ for (a) the test data set, (b) the basic neural ODE model, (c) the EDMD-structured neural ODE model, (d) EDMD with reprojection (e) the EDMD basis neural ODE, and (f) the EDMD direct neural ODE. The inset values give normalized Wasserstein distance $\overline{\mathcal{W}}_{2}$ between the distribution predicted by the model and the test data distribution. Note the logarithmic scale.
To assess the accuracy of the models in terms of reproducing the long-time statistics of the system, we simulate the models from the 100 initial conditions of the trajectories in the test dataset for 30,000 timesteps or until relaminarization is predicted. Shown in Fig. 8 are the joint probability density functions (PDFs) of the amplitudes $a_{1}$ and $a_{3}$, from the test data set and as predicted by each of the models. The standard EDMD model is omitted here, since it tends to rapidly diverge from the true trajectory, as shown in Fig. 7. From Fig. 8, it is apparent that each of the other models reproduces the statistics qualitatively well. To quantify the accuracy of these long-time statistics, we compute a Wasserstein distance (a.k.a. earth mover’s distance) to quantify the discrepancy between the predicted and true joint PDFs. This quantity measures the distance between distributions by solving an optimal tranport problem to determine the most efficient way to move mass between distributions. Given two discrete distributions, $\mu$ and $\nu$ (as obtained from a normalized histogram of the data using $n_{\mu}$ and $n_{\nu}$ bins, respectively), the Wasserstein-2 distance, $\mathcal{W}_{2}(\mu,\nu)$ between the distributions is given by
@@
where $\gamma\in\mathbb{R}^{n_{\mu}\times n_{\nu}}$ is a transport plan (i.e., $\gamma_{i,j}$ assigns the amount of mass to move from bin $\mu^{(i)}$ to bin $\nu^{(j)}$). We compute the Wasserstein distances using the python library, Python Optimal Transport [72], and we report the distances normalized by the distance between the distributions of the test and training datasets. That is, we report the normalized Wasserstein distance
@@
Figure 10 shows the $F$-score as a function of the prediction horizon normalized with Lyapunov time, $h/\tau_{L}$, for each of the models. From these results, it is clear that the continuous time neural ODE models outperform the discrete time and EDMD-DL-based approaches, more accurately forecasting extreme events for every prediction horizon. Among the EDMD-DL-based models, the EDMD-basis neural ODE outperforms EDMD-DL with reprojection and the EDMD direct neural ODE. The standard EDMD-DL prediction is poor for every prediction horizon, resulting in a much lower F-score than the other models. Also included for comparison in Fig. 10 are the $F$-score results from Racca and Magri [67], who developed a model using an echo state network, a type of recurrent neural network. Its predictions are based on long sequences of past states and are thus non-Markovian, while our model predictions are based only on the initial state. Nevertheless, their model only slightly outperforms the present models, especially for the case of the EDMD-structured neural ODE.

### v2

@@ III.4 Continuous-time models from EDMD
As noted in Sec. II, the Koopman formalism also gives rise to a continuous-time representation of the dynamics on the observable space, described by the infinitesimal generator, $\mathcal{L}$. Based on this, several previous works have shown that continuous-time representations can also be derived from EDMD or similar regression methods [47, 33, 48, 49, 50]. Using this continuous-time formulation, an analogous projection step can be performed, which directly yields a neural ODE representation of the state space dynamics, as illustrated in Fig. 2.
@@
In many implementations, it is assumed that time derivative data of the states $\dot{\mathbf{x}}$ is available (or can be obtained by finite-differencing). If this velocity data is available, then a matrix representation $\mathbf{L}$ of the Koopman generator, $\mathcal{L}$ can be obtained by minimizing the objective
@@
where $\dot{\Psi}(\mathbf{x})=\frac{\partial\Psi}{\partial\mathbf{x}}\cdot\dot{\mathbf{x}}$. This is the approach taken in [48, 49, 50].
Alternatively, one can obtain a matrix approximation of the generator from the EDMD approximation of the Koopman operator by applying the limit definition in Eq. 5 as
@@
though this approach will only be accurate for sufficiently small time interval, $\tau$. Similarly, the exponential relationship between the Koopman operator and the generator in Eq. 7 can be used to obtain a matrix approximation of $\mathcal{L}$ using a matrix logarithm as
@@
Once the matrix representation of the generator is obtained, it can be used to predict the dynamics of the state, $\mathbf{x}$, as this matrix acts on the dictionary elements to return their time derivative. Specifically, an initial state, $\mathbf{x}_{0}$, can be lifted to obtain an initial condition for the lifted state, $\Psi_{0}=\Psi(\mathbf{x}_{0})$, and then a trajectory in this lifted state can be obtained by integrating the equation
@@
Integrating this yields a trajectory in the lifted state, $\Psi$, from which we can recover the state predictions by mapping back to the state space, $\mathbf{x}=\mathcal{P}\Psi$, as in Sec. III.3. Note that Eq. 22 is a linear ODE in $\Psi$ and is essentially a continuous time version of the linear time evolution using the Koopman operator (without projection) in Eqs. 13, 16.
Analogously to the EDMD formulation, where a nonlinear discrete-time evolution equation on the state space was defined using the EDMD operator, $\mathbf{K}_{\tau}$, by projecting back to the state space on each timestep (see Eq. 17), a nonlinear continuous time evolution equation on the state space can also be defined using the EDMD generator, $\mathbf{L}$. This is done by lifting the state $\Psi(\mathbf{x})$, computing the time derivative of the feature vector, and then projecting the time derivative back to the state space. This can be seen by simply taking the time derivative of the projection back to the state space $\mathbf{x}=\mathcal{P}\Psi(\mathbf{x})$ as
@@
This form is only necessary if the mapping $\mathcal{P}$ back to the state space is nonlinear (such as in Koopman autoencoder formulations). If $\mathcal{P}$ is taken to be a linear mapping, $\mathcal{P}\Psi=\mathbf{P}\Psi$, such as when the state is explicitly included in the dictionary (see Sec. III.3, Eq. 14), then $\frac{\partial\mathcal{P}}{\partial\Psi}=\mathbf{P}$ and Eq. 23 can be simplified to
@@
In the case of EDMD-DL, where the feature vector $\Psi(\mathbf{x})$ is parametrized by a neural network, $\Psi(\mathbf{x};\theta_{\Psi})$ and optimized jointly with $\mathbf{K}_{\tau}$, this implies that the EDMD-DL training procedure directly defines a neural ODE of the form
@@
In summary, in both discrete time and continuous time, performing a projection back to the state space in the EDMD time evolution leads to a nonlinear dynamical system in the state space with a natural structure for neural networks. Specifically, the state is first expanded nonlinearly into a high dimensional feature space, followed by a linear mapping which represents the dynamics as a linear combination of these features. This observation motivates the development of several methods that combine different aspects of EDMD and neural ODE based models. In the following section, we show that adding the projection step to the EDMD-DL time evolution leads to much better predictive performance than the linear time evolution strategy used in standard EDMD-DL. Additionally, the models using this projection achieve a level of accuracy in predicting extreme events comparable to the non-Markovian approach of Racca and Magri [51].
Table 1: Summary of model structures based on neural ODEs and EDMD-DL.
@@
EDMD structured neural ODE $\dot{\mathbf{x}}=\bm{A}\Psi(\mathbf{x};\theta_{\Psi})$ $\theta_{\Psi}$ and matrix $\bm{A}=\mathbf{P}\mathbf{L}^{\mathrm{T}}$ optimized
@@
EDMD basis neural ODE $\dot{\mathbf{x}}=\bm{A}\Psi(\mathbf{x};\theta_{\Psi})$ $\theta_{\Psi}$ optimized by EDMD-DL (Eq. 15)
$\mathbf{x}(0)=\mathbf{x}_{0}$ $\bm{A}=\mathbf{P}\mathbf{L}^{\mathrm{T}}$ optimized as neural ODE (Eq. 9)
@@
IV Performance comparison
The discussion in Sec. III of the parallels between data-driven state-space modeling techniques and the projected Koopman approach to predicting the state evolution motivates a comparison between the numerical methods arising from each. In particular, we will compare two neural ODE models arising directly from the state space view and four neural network models arising from EDMD approximations of the Koopman operator (with learned dictionaries). These model variations are described in Sec. IV.1 below and summarized in Table 1. Then, in Sections IV.2 and IV.3 we present a performance comparison on two chaotic systems: the Lorenz system and a low-dimensional model of a turbulent shear flow. These examples demonstrate that while the standard, linear EDMD-DL fails in both cases, EDMD-DL with projection performs equivalently to the neural ODE models, which perform quite well across several metrics. This further supports our point that adding the projection step to the EDMD-DL time evolution makes the EDMD-DL model equivalent to a neural ODE, as it results in a nonlinear, neural network representation of the dynamics on the state space, rather than a linear representation of the dynamics on the space of observable functions as is typically associated with Koopman-based approaches.
While the examples presented here are low-dimensional, previous works have shown that neural ODE models [41, 52, 53, 54] and Koopman-based models with state-space projection [16, 55, 46] can be applied to higher dimensional systems, including chaotic PDEs such as the Kuramoto-Sivashinsky system, Kolmogorov flow, and turbulent Couette and pipe flows. For such high-dimensional systems, these studies paired dynamical models with autoencoder-based dimension reduction, learning the dynamics of a latent representation given by an autoencoder. In Ref. 41, it was shown that developing neural ODE models on such reduced representations can lead to better predictive performance than training a neural ODE directly on the higher dimensional data. Since the purpose of the numerical examples here is to demonstrate the equivalence of EDMD-DL with a state space projection to a neural ODE, we have chosen examples with complex, chaotic dynamics, but which are still sufficiently low dimensional so that additional dimension reduction steps are not needed. This allows us to assess the predictive capabilities of the dynamical models without introducing additional errors due to dimension reduction.
IV.1 Model structures
@@
The first EDMD approach that we consider is the standard EDMD-DL method, where a dictionary parametrized by a neural network and the corresponding Koopman operator matrix representation are jointly learned by optimizing the objective in Eq. 15. Then to model the time-evolution, the initial state is lifted and propagated linearly in the lifted space using the Koopman matrix (see Eq. 16). We also consider the EDMD variation with projection, where the same dictionary and Koopman matrix are used, but where on each timestep, the state is lifted, evolved forward in time, and then reprojected (see. Eq. 17). In each EDMD-DL implementation, we constrain the dictionary to be of the form in Eq. 14, where the state is explicitly included as a nontrainable set of dictionary elements.
@@
We consider two variations of continuous-time models derived from EDMD. In the first, we use the dictionary obtained from EDMD-DL as the neural network in a neural ODE. That is, the dictionary is optimized, as in EDMD-DL, to minimize Eq. 15, and then the linear layer that maps to the output is optimized through the neural ODE training procedure. So, the model has the structure of Eq. 25, where the dictionary $\Psi(\mathbf{x};\theta_{\Psi})$ is determined by EDMD-DL and the linear mapping $\bm{A}\equiv\mathbf{P}\mathbf{L}^{\mathrm{T}}$ is trained as with the neural ODE loss in Eq. 9. We refer to this variation as an EDMD basis neural ODE. Lastly, we consider a neural ODE model that is obtained directly from EDMD-DL, without a separate training procedure. That is, the matrix representation of the generator $\mathbf{L}$ is obtained from the EDMD-DL Koopman matrix $\mathbf{K}_{\tau}$ by a matrix logarithm, as in Eq. 21. We refer to this variation as an EDMD direct neural ODE.
@@ IV.2 Lorenz system
Finally, while Figs. 3 and 5 analyzed the short-time performance of these models, we demonstrate the long-time performance of the models by simulating each model for $100$ time units from a common initial condition. The resulting trajectory from each model is shown in Fig. 4. We see that each of the models considered (excluding the standard EDMD-DL approach) are able to capture the dynamics and reproduce the expected butterfly shape of the Lorenz attractor.
@@ IV.3 Low dimensional model of turbulent shear flow
The domain is wall-bounded at $y=\pm 1$ and periodic in the streamwise and spanwise directions ($x$ and $z$), with lengths $L_{x}=4\pi$ and $L_{z}=2\pi$, respectively; and we consider a channel Reynolds number of $\mathrm{Re}=400$, following Refs. 63, 51, 15, 46. The Lyapunov timescale associated with these parameter values is $\tau_{L}\approx 61$ [51]. The nine mode model is constructed as a Galerkin projection of the Navier-Stokes equations onto nine Fourier modes, as described in Ref. 56. That is, the fluid velocity field $\mathbf{u}(\mathbf{x},t)$ is approximated as a superposition of the modes, $\mathbf{u}_{i}(\mathbf{x})$, with time-varying amplitudes $a_{i}(t)$
@@
Performing the Galerkin projection yields a set of nine ODEs for the amplitudes. These ODEs, as well as the Fourier modes for the model, are given explicitly in Ref. 56. The first mode of the model, $\mathbf{u}_{1}(\mathbf{x})$, represents the base flow profile, so the laminar flow state is associated with the amplitudes $\mathbf{a}=[1,0,0,0,0,0,0,0,0]^{\mathrm{T}}$. The flow can be characterized using the energy $E(t)=\sum_{i=1}^{9}a_{i}^{2}(t)$. In the turbulent region of the state space, the system exhibits extreme, high energy quasi-laminarization events, while the laminar state corresponds to the steady value of $E=1$. A typical time-series of the energy for this system is shown in Fig. 7
@@
A training dataset is constructed by numerically integrating the system from 100 random initial conditions, $\mathbf{a}(0)\sim\mathcal{U}([-0.1,0.1]^{9})$ for 30,000 time units or until laminarization occurs with data stored at time intervals of $\tau=0.5$. The first $500$ time units of each trajectory are discarded as transience. These trajectories have a mean lifetime of approximately $1.1\times 10^{4}$ (or $180\,\tau_{L}$) and the resulting dataset contains approximately $2.1\times 10^{6}$ snapshots. A test dataset is generated using an identical procedure, using another set of randomly selected initial conditions. All models are trained for 10 epochs with batch size of 40 ($5.25\times 10^{5}$ iterations).
@@
Figure 8: Joint probability densities of amplitudes $a_{1}$ and $a_{3}$ for (a) the test data set, (b) the basic neural ODE model, (c) the EDMD-structured neural ODE model, (d) EDMD with projection (e) the EDMD basis neural ODE, and (f) the EDMD direct neural ODE. The inset values give normalized Wasserstein distance $\overline{\mathcal{W}}_{2}$ between the distribution predicted by the model and the test data distribution. Note the logarithmic scale.
To assess the accuracy of the models in terms of reproducing the long-time statistics of the system, we simulate the models from the 100 initial conditions of the trajectories in the test dataset for 30,000 timesteps or until relaminarization is predicted. Shown in Fig. 8 are the joint probability density functions (PDFs) of the amplitudes $a_{1}$ and $a_{3}$, from this test data set and as predicted by each of the models. The standard EDMD model is omitted here, since it tends to rapidly diverge from the true trajectory, as shown in Fig. 6. From Fig. 8, it is apparent that each of the other models reproduces the statistics qualitatively well. To quantify the accuracy of these long-time statistics, we compute a Wasserstein distance (a.k.a. earth mover’s distance) to quantify the discrepancy between the predicted and true joint PDFs. For this, we consider the empirical distribution of the test distribution as the true joint PDF. The Wasserstein distance measures the distance between distributions by solving an optimal tranport problem to determine the most efficient way to move mass between distributions. Given two discrete distributions, $\mu$ and $\nu$ (as obtained from a normalized histogram of the data using $n_{\mu}$ and $n_{\nu}$ bins, respectively), the Wasserstein-2 distance, $\mathcal{W}_{2}(\mu,\nu)$ between the distributions is given by
@@
where $\gamma\in\mathbb{R}^{n_{\mu}\times n_{\nu}}$ is a transport plan (i.e., $\gamma_{i,j}$ assigns the amount of mass to move from bin $\mu^{(i)}$ to bin $\nu^{(j)}$). Other distances for quanitatively comparing distributions exist, such as the Kullback-Leibler divergence or a norm of the error, but these distances are typically more sensitive to small shifts of the distribution or the discretization of the distribution than the Wasserstein distance. For a more detailed discussion of the use of the Wasserstein distance and a comparison of its properties to those of other statistical divergences commonly used in data science, we refer to Peyré and Cuturi [64]. We compute the Wasserstein distances using the python library, Python Optimal Transport [65], and we report the distances normalized by the distance between the distributions of the test and training datasets. That is, we report the normalized Wasserstein distance
@@
Figure 10 shows the $F$-score as a function of the prediction horizon normalized with Lyapunov time, $h/\tau_{L}$, for each of the models. From these results, it is clear that the continuous time neural ODE models outperform the discrete time and EDMD-DL-based approaches, more accurately forecasting extreme events for every prediction horizon. Among the EDMD-DL-based models, the EDMD-basis neural ODE outperforms EDMD-DL with projection and the EDMD direct neural ODE. The standard EDMD-DL prediction is poor for every prediction horizon, resulting in a much lower $F$-score than the other models. Also included for comparison in Fig. 10 are the $F$-score results from Racca and Magri [51], who developed a model using an echo state network, a type of recurrent neural network. Its predictions are based on long sequences of past states and are thus non-Markovian, while our model predictions are based only on the initial state. Nevertheless, their model only slightly outperforms the present models, especially for the case of the EDMD-structured neural ODE.
While these $F$-score results results demonstrate the short-time prediction of extreme events, next we consider the long-time statistical accuracy of extreme-event prediction for each of these models. For this, we compute the ergodic average of an extreme event indicator function. This indicator function is defined as
$I_{EE}(\mathbf{a})=\begin{cases}1\qquad E(\mathbf{a})>0.2\\ 0\qquad E(\mathbf{a})\leq 0.2\end{cases}$ (32)
where $E(\mathbf{a})$ is the kinetic energy, as defined above. The ergodic average of this indicator essentially represents the fraction of time that trajectories spend in the high kinetic energy, near-laminar region of the state space. The ergodic average of the extreme event indicator function for each of the models is computed over the trajectory datasets used in Fig. 8, which consist of simulations from each model from 100 initial conditions for 30,000 timesteps or until relaminarization is predicted. The resulting ergodic averages are given in Table 2.
Table 2: Ergodic average of extreme-event indicator function.
Model $\langle I_{EE}\rangle$
Test data  0.130
Basic neural ODE 0.125
EDMD structured neural ODE 0.131
EDMD-DL with projection 0.128
EDMD basis neural ODE 0.142
EDMD direct neural ODE 0.131
This result further demonstrates that EDMD-DL with projection and each of the neural ODE models perform comparably, yielding an ergodic average of this indicator that is close to the value when calculated on the test data. This indicates that these models successfully capture the long-time statistics of extreme events.